# Hierarchical Object-Oriented POMDP Planning for Object Rearrangement

### v2

@@ 1 Function HOO_POMDP_Rearrangement():
    • Done - This action receives a reward of $50$ if all objects have been placed at their goal location and $-50$ otherwise.
Abstract OOPOMDP Planner: Given a task defined as an abstract OOPOMDP and an initial abstract state, the OOPOMDP planner uses partially observable UCT (POUCT) Silver & Veness (2010) to search through the space of abstract actions to find the best sub-goal. POUCT is an extension of the UCT algorithm Kocsis & Szepesvári (2006) to partially observable settings. The search tree in POUCT is over the histories instead of states. A history is a sequence of actions and observations $h_{t}=(a_{1},z_{1},\ldots,a_{t},z_{t})$. Similar to the UCT algorithm, the POUCT maintains a tree, where each node corresponds to a history $T(h)$, and for each node, a count variable ($N(h)$) and a value variable ($v(h)$) are stored, representing the number of times this history has been visited and the expected value of history $h$, which is estimated by the expected return of all simulations starting at $h$. The algorithm samples a state from the belief space $b$ for the current history, and if the tree already contains all the children for the current node, then the action with the best value is selected using the equation $V(ha)=V(ha)+c\sqrt{\frac{\log N(h)}{N(ha)}}$ to compute the value for all actions. If the tree does not contain all children, then a rollout policy is used to select actions for simulations, and the tree is updated with information from the simulations. Once that is done, the best action is selected and returned. The full algorithm is in Algorithm 3 in the Appendix. The rollout policy is a random policy that picks randomly amongst available actions. The Move and the Rotate actions are initialized by the values in the abstract state for each object. Recall that each object has the following information in its abstract state $s_{i}=(loc_{i},\ pick_{i},\ place_{locs},\ is\_held,\ at\_goal)$. A separate $\textit{Move}_{AB}$ is initialized with $A=agent\_pos$ and $B=$ all pick locations defined for all objects. $\textit{Rotate}_{angle}$ for all objects, less than 2m from the agent, the angle is computed based on the agent’s required orientation to view the object from its current position. PickPlace - is defined for each object where the agent is less than 2m away from that particular object, for all locations in the $place_{locs}$ as $goal_{loc}$ initializing a set of PickPlace action for each object.
Policy Executor: Each sub-goal/action in the abstract OO-POMDP corresponds to a policy. When the planner outputs a sub-goal, the information in the given sub-goal is used to initialize the low-level policy. The output from the low-level policy is a sequence of low-level actions.
@@
  • Sub-goal $\textit{Rotate}_{angle}$ gives the Rotate policy the final angle to be at, which is used as the final state the $A^{*}$ system must reach and outputs a sequence of rotate actions.
  • Sub-goal $\textit{PickPlace}_{Object_{i}-goal_{loc}}$ provides the object to interact with and which location to place it at. The policy takes this information as input and outputs a sequence of actions consisting of Pick, Place, and navigation actions that enable it to pick the object, move to the given destination, and place the object at that location. The PickPlace consists of 3 separate components a) An RL model trained to pick an object, b) the $A^{*}$ navigation model to go its destination c) An RL model trained to place the object when the agent is near the goal. All 3 of these run sequentially and make up the PickPlace Policy. It is designed this way to improve modularity and reduce the complexity of each part. Both the pick model and place model are trained using the PPO algorithm. The action space is $\{A_{s}-\textit{Place}\}$ for pick model and $\{A_{s}-\textit{Pick}\}$ for the place model. The training process for the pick model involves randomly positioning the target object within a specified proximity to the agent. The goal is to pick a selected object successfully. For the place model, the training methodology follows a similar approach, with the key distinction being the absence of object detection requirements, as the agent begins each scenario already holding the object. In all training instances for the place model, the initial state consists of the agent holding an object, and the task involves depositing the object at a predetermined location.
Perception: Once the agent executes the low-level action, it receives an RGB and Depth image. An object detector is used to detect objects in the RGB image, and the depth map is used to get their $3D$ location in the world. This is used to generate the object-oriented observation $z=(z_{1},\ldots,z_{n})$.
Belief Update: Algorithm 2 presents the belief update function for our HOO-POMDP. The UpdateBelief function takes as input the current belief state $b$, the performed action $a$, and the received observation $z$. For each object $i$ in the belief state and each possible state $s_{ij}$ ($j=1,\ldots,L$, where $L$ is the set of all its possible locations in the 2D Map) of that object, the algorithm updates the belief based on the action type. For navigation actions, it applies a probabilistic update using the observation model $p(z_{i}|s_{ij})$ (Line 2). For ‘place’ actions, it sets the belief to 1 if the object’s state matches its desired goal location $g_{i}$, and 0 otherwise (Line 2). For ‘pick’ actions, it assigns a belief of 1 if the object’s state corresponds to the agent’s location and 0 otherwise (Line 2).
1
2Function UpdateBelief(beliefState, action, observation):
3       $b\leftarrow$ beliefState ; $z\leftarrow$ observation;
4       for each object $i$ in $b$ do
5             for each possible state $s_{ij}$ of object $i$ do
6                   if action is navigation then
7                         $b^{\prime}_{i}(s_{ij})\leftarrow\eta p(z_{i}|s_{ij})b_{i}(s_{ij})$
8
9                  else if action is place then
[…]
11                  else if action is pick then
[…]
13      return $b^{\prime}$;
14
Algorithm 2 HOO-POMDP Belief Update
@@ 3       $b\leftarrow$ beliefState ; $z\leftarrow$ observation;
The value for $is\_held$ comes from the previous low-level action and previous state. If the previous state had $is\_held$ as false and low-level action was to pick the object of interest, $is\_held$ is set to true. If the previous action was not a pick or a pick action for a different object, then the variable remains unchanged. If the previous action was $place$ and $is\_held$ is true, then it is set to False.
The value for $at\_goal$ is copied from the previous state if the last low-level action was not the place action. If it was, and if $is\_held$ was true in the previous state, then $at\_goal$ is set to true.
The value for $place_{locs}$ are sampled from the object goal location and three nearby receptacles as alternate goal locations for the object. For each of these goal locations, a location from where the object can be placed is sampled.
The location $pick_{i}$ is sampled based on the belief distribution of where the object could be and is the location from which the object can be picked. If the distribution over location is spread out, we sample multiple locations (by ensuring each sampled location is far from the other sampled locations for the same object). For both the locations in $place_{locs}$ and for the location $pick_{i}$, we then check if they are reachable. If they are not, those locations are discarded.
This sampling method enables our system to handle scenarios involving blocked goals, object swaps, and obstructed paths effectively. If an object’s path is blocked, the planner will receive information indicating that there is no accessible location from which to pick up the object, necessitating the relocation of other objects first. When placing objects, we provide alternative receptacle locations. This approach allows us to move an object to another location if its goal position is blocked, thereby freeing up its current location. This strategy addresses both blocked goal and swap scenarios. Furthermore, this sampling process enhances our system’s extensibility. We can incorporate additional constraints based on new object properties. For example, if opening an object requires interaction from a distance, the sampler can ensure that the sampled location is sufficiently far to enable successful opening.
After creating this abstract state, it is sent to the abstract planner, and the cycle starts again.
@@ 5.2 Metrics
  • Success Rate (SR): 1 if all objects have been moved to the correct goal locations, 0 otherwise.
  • Object Success Rate (OSR): (Total Objects successfully moved)/(Total objects to move) - this metric captures the proportion of objects moved to the correct goal location.
  • Total Actions taken (TA): A measure of the efficiency of the system in terms of the total number of actions taken. We present the average number(rounded up) of actions taken during successful runs where the scene was fully rearranged.
5.3 Baselines definition
  • Perfect Knowledge (PK): In this baseline, we will start with all the information about the world. That is, we know the initial locations of all the objects. This is the upper limit of the system’s performance as there is no uncertainty to manage.
  • Perfect Detector with partial observability (PD): In this baseline, we will be solving our multi-object rearrangement problem with a perfect detector (objects in the visual field are detected with 100% probability). The main challenge is to find all objects and move them around efficiently.
  • OURS (Imperfect Detector with partial observability): In this setting, we will solve the multi-object rearrangement challenge where the agent is expected to handle perception uncertainty (the detector fails to detect objects in the visual field) along with object’s starting position uncertainty.
  • OURS-HP (ablation): In this setting, we remove the hierarchical planning and use the POMDP planner to output low-level actions directly.
Experimental setup: Each experimental setting was evaluated across 100 distinct rearrangement configurations. For the RoomR dataset, we utilized 25 different room setups, with four rearrangement configurations per setup. For the other datasets, we employed 100 unique room configurations, each with one rearrangement configuration. In the blocked goal and swap settings, at least one object’s goal location was obstructed by another object. The swap case required the exchange of positions for at least one pair of objects to complete the rearrangement task. In the blocked path scenario, all scenes contained a minimum of one object that needed to be moved to its goal position from its initial location to enable the rearrangement of other objects.
@@ 6 Results and Discussion
Dataset Objs #BG #Sw #BP #Rm #V PK PD Ours OURS-HP
@@
RoomR 5 1 0 0 1 3-4 63 88 176 62 87 189 49 71 211 13 33 302
2 1 1 1 3-4 51 77 210 49 76 238 39 61 289 8 27 392
Proc 5 1 0 0 2 2-3 60 82 203 60 81 269 46 68 352 9 29 410
2 1 1 2 2-3 47 71 246 42 63 311 31 53 398 4 19 565
Multi RoomR 10 1 1 0 2 2-3 41 77 457 40 79 529 32 65 710 5 25 1029
2 1 1 2 2-3 33 69 489 29 67 587 21 49 789 2 19 1092
10 2 1 0 3-4 1-2 39 75 726 37 74 834 30 62 1189 3 16 1408
2 1 1 3-4 1-2 32 69 789 31 72 985 18 44 1321 1 7 1549
15 1 1 0 3-4 2-3 32 78 895 30 74 921 22 59 1228 NC NC NC
2 1 1 3-4 2-3 29 71 988 25 69 965 14 41 1416 NC NC NC
20 2 1 0 3-4 2-4 27 75 1168 27 74 1197 17 55 1621 NC NC NC
2 1 1 3-4 2-4 22 70 1307 20 68 1336 10 36 1786 NC NC NC
Table 1: Comparison between our methods with different levels of information and difficulty of problems and ablation on hierarchy. The difficulty is represented in terms of the following: a) #BG - number of blocked goal locations, b) #Swap Number of objects that need to be swapped, c) #BP: Number of objects blocking the path that need to be moved out of the way, d) #Rm: Number of rooms in the environment, e) #V: Number of objects initially visible. NC: Not computed.
Methods comparison: The perfect knowledge system is the best performance our agent can have - there is no uncertainty in the world, and hence, it boils down to a path-traversal length optimization problem. We can see from Table 1 that the results of the perfect detector setting and the perfect knowledge setting are similar in the overall scene success rate as well as the object success rate. This shows that the PPOMDP planner is able to explore and find all the objects perfectly. The failures in both cases are due to the low-level policies failing to pick/place objects. The difference between the methods can be seen in the steps taken. The perfect detector system has to explore and rearrange and hence ends up taking more steps than the perfect knowledge system. The third set of results are for our system with an imperfect detector and partial observability. We can see that the detector failure causes our success rate to fall by a fair amount - this is because if an object is not detected more than once at its location, the planner’s belief update prevents it from going to that region to pick that object up. Since our detector only has a 50-60% success rate (different for different classes of objects), it misses a fair subset of the objects, and hence the drop. As a result of needing to explore more due to detector failure, the number of steps taken is greater than that of the other methods. The difference between OURS and OURS-HP clearly shows the importance of hierarchy and abstraction. The performance drop is significant across all settings for OURS-HP.
Comparison Across Datasets: We can see that the results of the RoomR dataset of a single room and the Proc dataset with multiple rooms are similar (in scene success rate and object success rate). This indicates that our method effectively manages the exploration of multiple rooms as well as a reduction in the number of objects visible initially. Initial object visibility decreases when scaling from 2 to 4 rooms in MultiRoomR with ten objects, yet success rates remain relatively stable. The performance drops in our custom dataset with a larger number of objects. In particular, the object success rate drops slightly, but the scene success rate drops significantly. This is because, for a scene to succeed, we need all objects to be rearranged correctly, so even if 9/10 objects are rearranged correctly, it is still considered a failure. Hence, there is a bigger reduction in scene success rate.
Across Different Challenges: From the different rows for each dataset (the difference being the existence of blocked paths), we can see that the results for the blocked path version of the problem are lower. This is because the low-level policy is not as good at picking the objects on the floor as it is at picking the objects from the top of other objects. Also, the number of PickPlace actions increases as we tackle more complex problems such as swaps.
Error Analysis: The majority of failures of our system are due to low-level policy failures - the pick or the place action fails due to the imperfections of the low-level RL policy. The other type of failure in our system is due to belief estimation errors caused by detector failures - if a detector is looking at a certain location and it fails to detect an object multiple times (due to partial occlusion or high distance), then our belief about the object being in that region reduces to a large extent. In that case, we are unlikely to come back to that part of the house. If a false positive happens, an agent might pick up the wrong object and place it and believe it has placed the correct object - leading to object and scene failure. These are the main causes of our failures in imperfect detector settings.
Comparison to existing baselines: It is important to note that our system addresses a variant of the multi-object rearrangement problem that differs in key aspects from those tackled by existing baselines. Unfortunately, this means that direct comparisons are not meaningful or informative. The primary distinction lies in the prior knowledge available to our system: we are given information about the classes of objects to be moved, whereas other systems Mirakhor et al. (2024a) (Mirakhor et al., 2024b), (Gadre et al., 2022) operate without this advantage. The motivation for it is that when this information is not provided, agents must perform a walkthrough phase for each new goal configuration to identify movable objects. In contrast, our formulation requires only a single initial walkthrough to map stationary objects in the environment. Subsequently, our system can efficiently handle multiple goal configurations without additional walkthroughs.
@@
However, it is worth highlighting that our problem formulation introduces its own set of challenges. In particular, while existing systems report initial visibility of approximately 60% ((Mirakhor et al., 2024b), table 1) of target objects at the outset of their tasks, only about 20% of the objects are initially visible in our problem settings, necessitating more extensive and strategic exploration. This reduced initial visibility significantly increases the complexity of our task in terms of efficient exploration and belief management. It underscores the importance and effectiveness of our hierarchical planning approach in handling partial observability and perception uncertainty. The other key difference is we use low-level policies for navigation and manipulation, whereas other works Mirakhor et al. (2024b) and Mirakhor et al. (2024a) assume perfect navigation and manipulation capabilities.
While the difficulties faced by our system and existing baselines are not directly comparable, we believe our results demonstrate the efficacy of our approach in solving a challenging and practically relevant variant of the multi-object rearrangement problem. The performance across various metrics and scenarios, as detailed in the preceding sections, showcases the robustness and scalability of our HOO-POMDP framework in environments with high uncertainty and limited initial information.
@@
Limitations: Our system makes an object independence assumption - that the state and observation of one object are not dependent on the state of other objects. However, this assumption doesn’t always hold in cluttered environments. In such cases, observing, picking up, or placing an object might be affected by the positions of nearby objects. As a result, our object-oriented belief update system would need to be modified to handle these object interactions in more complex scenarios. HOO-POMDP cannot handle an unknown class of objects. It could potentially be handled by categorizing all of the known object types into a single ‘unknown’ class. The difficult part, however, is to plan to find an empty space to move the unknown object to. In the worst case, this could lead to complicated packing problems that are NP-hard.

### v3

@@ 4 Hierarchical Object Oriented POMDP (HOO-POMDP) Framework
    • Done - This action receives a reward of $50$ if all objects have been placed at their goal location and 50-50 otherwise.
Abstract OOPOMDP Planner: Given a task defined as an abstract OOPOMDP and an initial abstract state, we use partially observable UCT (PO-UCT) (Silver and Veness, 2010) to search through the space of abstract actions to find the best sub-goal. POUCT extends the UCT algorithm Kocsis and Szepesvári (2006) to partially observable settings, building a search tree over histories rather than states, where a history is a sequence of actions and observations $h_{t}=(a_{1},z_{1},\ldots,a_{t},z_{t})$. For each history node $T(h)$, the algorithm maintains a count variable $N(h)$ and a value variable $v(h)$ representing visit frequency and expected value. The algorithm samples a state from the belief space $b$ and selects the action with the best value using $V(ha)=V(ha)+c\sqrt{\frac{\log N(h)}{N(ha)}}$ when all child nodes exist. Otherwise, it uses a random rollout policy for simulations before updating the tree and selecting the best action. The full algorithm appears in Algorithm 3 in the Appendix.
In our HOO-POMDP framework, abstract actions are initialized based on the abstract state for each object $s_{i}=(loc_{i},\ pick_{i},\ place_{locs},\ is\_held,\ at\_goal)$. A separate $\textit{Move}_{AB}$ is initialized with A=agent_posA=agent\_pos and $B=$ all pick locations defined for all objects. $\textit{Rotate}_{angle}$ - for all objects, less than 2m from the agent, the angle is computed based on the agent’s required orientation to view the object from its current position. PickPlace - is defined for each object where the agent is less than 2m away from that particular object, for all locations in the $place_{locs}$ as $goal_{loc}$ initializing a set of PickPlace actions for each object.
Low-Level Policy Executor: Each sub-goal (instantiated abstract action) output by the abstract planner corresponds to a policy. When the planner outputs a sub-goal, the information in the sub-goal is used to initialize the low-level policy. The output from the low-level policy is a sequence of low-level actions. We then execute the first low-level action in this sequence in the environment.
@@
  • Sub-goal $\textit{Rotate}_{angle}$ gives the Rotate policy the final angle to be at, which is used as the final state the $A^{*}$ system must reach. $A^{*}$ outputs a sequence of rotate actions.
  • Sub-goal $\textit{PickPlace}_{Object_{i}-goal_{loc}}$ provides the object to interact with and which location to place it at. The policy takes this information as input and outputs a sequence of actions consisting of Pick, Place, and navigation actions. The PickPlace consists of 3 separate components a) An RL model trained to pick an object, b) the $A^{*}$ navigation model to go its destination c) An RL model trained to place the object when the agent is near the goal. All 3 of these run sequentially and make up the PickPlace Policy. It is designed this way to improve modularity and reduce the complexity of each part. All details of the RL policy training are described in appendix A.5
Perception System: Once the agent executes the low-level action, it receives an RGB and Depth image. An detector is used to detect objects in the RGB image, and the depth map is used to get their $3D$ location in the world. This is used to generate the object-oriented observation $z=(z_{1},\ldots,z_{n})$.
Belief Update: Algorithm 2 presents the belief update function for our HOO-POMDP. The UpdateBelief function takes as input the current belief state $b$, the performed action $a$, and the received observation $z$ (Line 1.). For each object $i$ in the belief state and each possible state $s_{ij}$ ($j=1,\ldots,L$, where $L$ is the set of all its possible locations in $M^{2D}$) of that object, the algorithm updates the belief based on the action type. For navigation actions, it applies a probabilistic update using the observation model $p(z_{i}|s_{ij})$ (Line 5.). For ‘place’ actions, it sets the belief to 1 if the object’s state matches its desired goal location $g_{i}$, and 0 otherwise (Line 7.). For ‘pick’ actions, it assigns a belief of 1 if the object’s state corresponds to the agent’s location and 0 otherwise (Line 9.).
@@
The value for is_heldis\_held comes from the previous low-level action and previous state. If the previous state had is_heldis\_held as false and low-level action was to pick the object of interest, is_heldis\_held is set to true. If the previous action was not a pick or a pick action for a different object, then the variable remains unchanged. If the previous action was placeplace and is_heldis\_held is true, then it is set to False.
The value for at_goalat\_goal is copied from the previous state if the last low-level action was not the place action. If it was, and if is_heldis\_held was true in the previous state, then at_goalat\_goal is set to true.
The values for $place_{locs}$ are sampled from the object goal location and three nearby receptacles as alternate goal locations for the object. For each of these goal locations, a location from where the object can be placed is sampled.
The location $pick_{i}$ is sampled based on the belief distribution of where the object could be. It is the location from which the object can be picked. If the distribution over location is spread out, we sample multiple locations (by ensuring each sampled location is far from the other sampled locations for the same object). For both the locations in $place_{locs}$ and for the location $pick_{i}$, we then check if they are reachable. If they are not, those locations are discarded.
This sampling method enables our system to handle scenarios involving blocked goals, object swaps, and blocked paths effectively. If an object’s path is blocked, the planner will receive information indicating that there is no accessible location from which to pick up the object, necessitating the relocation of other objects first. When placing objects, we provide alternative receptacle locations. This approach allows us to move an object to another location if its goal position is blocked, thereby freeing up its current location. This strategy addresses both blocked goal and swap scenarios. Furthermore, this sampling process enhances our system’s extensibility. We can incorporate additional constraints based on new object properties. For example, if opening an object requires interaction from a distance, the sampler can ensure that the sampled location is sufficiently far to enable successful opening. After creating this abstract state, it is sent to the abstract planner, and the cycle starts again.
@@ 5.2 Metrics
  • Success Success (SS): 1 if all objects have been moved to the correct goal locations, 0 otherwise.
  • Object Success Rate (OS): (Total Objects successfully moved)/(Total objects to move) - this metric captures the proportion of objects moved to the correct goal location.
  • Total Actions taken (TA): The average number (rounded up) of actions taken during successful runs where the scene was fully rearranged, which is a measure of the efficiency of the system.
5.3 Methods and Baselines definition
  • HOOP (HOO-POMDP Planner): Our proposed solution. In this, we will solve the rearrangement challenge where the agent handles perception uncertainty (the detector fails to detect objects in the visual field) along with the object’s position uncertainty using the proposed HOO-POMDP planner.
  • Frontier Exploration + Hand-Coded Interaction (FHC) : This baseline employs a frontier exploration strategy that systematically explores the environment until objects that need rearrangement are discovered. It uses a confidence-based approach where an object is considered detected when the belief probability exceeds 70%. Once detected, the object is immediately rearranged and the system resumes exploration for the next undetected object. More details in appendix A.6
  • VRR: This is the model from Weihs et al. (2021) trains and RL agent using PPO (Schulman et al., 2017) and imitation learning with RGB images as input, builds a semantic map using Active Neural Slam (Chaplot et al., 2020) and outputs low-level actions directly.
  • MSS : System from Trabucco et al. (2022) first builds a 3D map of the whole world, then navigates to each object that needs to be moved and rearranges it.
  • HOOP-HP (ablation): In this ablation setting, we remove the hierarchical planning and use the POMDP planner to output low-level actions directly.
  • Perfect Knowledge (PK): In this oracle setting of our system, we will start with all the information about the world. That is, we know the initial locations of all the objects. This is the upper limit of the system’s performance as there is no uncertainty to manage.
  • Perfect Detector with partial observability (PD): In this oracle setting of our system, we solve our multi-object rearrangement problem with a perfect detector (objects in the visual field are detected with 100% probability). The challenge is to find all objects and move them around efficiently.
Experimental setup: Each experimental setting was evaluated across 100 distinct rearrangement configurations. For the RoomR dataset, we utilized 25 different room setups, with four rearrangement configurations per setup. For the other datasets, we employed 100 unique room configurations, each with one rearrangement configuration. In the blocked path scenario, all scenes contained a minimum of one object that needed to be moved to its goal position from its initial location to enable the rearrangement of other objects. Results for blocked goal and swap cases in table 3 in appendix.
@@ 6 Results and Discussion
Table 1: Comparison between all methods. The difficulty is represented in terms of the following: a) #BP: Number of objects blocking the path that need to be moved out of the way, b) #Rm: Number of rooms in the environment, c) #V: Number of objects initially visible. NA: Not applicable as no scene was fully rearranged. NC - Not computable as the method cannot handle blocked paths
Methods Ablation Oracle Settings
Dataset Objs #BP #Rm #V HOOP FHC VRR MSS HOOP-HP PK PD
@@
RoomR 5 0 1 3-4 49 71 211 38 58 269 7 31 256 21 44 267 13 33 302 63 88 176 62 87 189
Proc 5 0 2 2-3 46 68 352 32 61 411 2 19 382 14 29 395 9 29 410 60 82 203 60 81 269
Multi RoomR 10 0 2 2-3 32 65 710 20 44 931 0 13 NA 8 25 920 5 25 1029 41 78 457 40 78 529
10 1 2 2-3 21 49 789 12 38 993 0 9 NA NC NC NC 2 19 1092 33 69 489 29 67 587
10 0 3-4 1-2 30 62 1189 19 34 1345 0 8 NA 0 14 NA 3 16 1408 39 75 726 37 74 834
10 1 3-4 1-2 18 44 1321 9 26 1490 0 5 NA NC NC NC 1 7 1549 32 70 789 31 70 985
15 0 3-4 2-3 22 59 1228 12 31 1605 0 9 NA 0 11 NA 0 5 NA 32 78 895 30 74 921
15 1 3-4 2-3 14 41 1416 7 23 1886 0 5 NA NC NC NC 0 6 NA 29 71 988 25 69 965
20 0 3-4 2-4 17 55 1621 0 18 NA 0 6 NA 0 9 NA 0 5 NA 27 75 1168 27 74 1197
20 1 3-4 2-4 10 36 1786 0 11 NA 0 4 NA NC NC NC 0 4 NA 22 70 1307 20 68 1336
Methods comparison:
PK represents our agent’s upper performance bound with no uncertainty. As shown in Table 1, the PD setting achieves comparable scene success and object success rates to PK, demonstrating our POMDP planner effectively handles partial observability. The primary difference appears in action count—PD requires more steps than PK due to PD actually needing to look for the objects and plan with this uncertainty. HOOP system with imperfect detection and partial observability shows reduced success rates as the planner must account for detection failures. Despite only 50-60% detection success across object classes, HOOP still solves many problems comparable to PD, demonstrating its effectiveness in handling detector failures. The increased exploration necessitated by detector failures results in more steps versus other methods. The substantial performance gap between HOOP and HOOP-HP demonstrates the critical importance of our hierarchical abstraction approach.
Comparison to baselines : Our system significantly outperforms all baselines, demonstrating the clear advantages of a principled planning approach. While VRR’s pure reinforcement learning strategy struggles to scale beyond simple environments, failing dramatically as complexity increases, both hand-coded methods (FHC and MSS) also show substantial limitations. FHC not only struggles with complex spatial reasoning but also fails to handle object detection failures effectively, as it lacks a belief-based framework to account for perceptual uncertainty. MSS’s rigid explore-then-rearrange strategy prevents it from addressing blocked paths and causes inefficient execution even in simpler scenarios. These results highlight that planning-based approaches provide essential flexibility and robustness for real-world rearrangement tasks where uncertainty and complexity are inevitable, outperforming both end-to-end learning and hand-engineered heuristic methods.
@@
It is important to note that our system addresses a variant of the multi-object rearrangement problem that differs in key aspects from those tackled by existing baselines. The primary distinction is that we are given information about the classes of objects to be moved, whereas other systems, VRR (Weihs et al., 2021) and MSS (Trabucco et al., 2022)operate without this knowledge. On the other hand, our problem formulation introduces its own set of challenges. In particular, while existing systems report initial visibility of approximately 60% ((Mirakhor et al., 2024b), Table 1) of target objects at the outset of their tasks, only about 20% of the objects are initially visible in our problem settings in MultiRoomR, necessitating more extensive and strategic exploration. This reduced initial visibility significantly increases the complexity of our task in terms of efficient exploration and belief management, and underscores the effectiveness of our approach.
Comparison across datasets : Our system maintains relatively consistent performance across datasets when controlling for room and object count. The RoomR (single room) and Proc (two rooms) datasets show similar success rates despite differing room counts. As we scale to MultiRoomR with more objects, object success rate decreases gradually while scene success drops more sharply since it requires all objects to be successfully rearranged.
Performance across challenges : The blocked path variants consistently show lower success rates due to the increased complexity and limitations in our low-level manipulation policy, which struggles more with floor objects than tabletop ones. The required PickPlace actions increase for blocked path scenarios as the agent must execute intermediate movements to clear pathways.
Error analysis : Most failures stem from two factors: low-level policy limitations in pick/place actions and belief estimation errors from detector failures. When the detector misses an object multiple times, our belief about its presence decreases, making the planner unlikely to revisit that area. False positives can also cause manipulation of incorrect objects. Despite these challenges, our approach significantly outperforms baselines across diverse scenarios.
@@
Limitations: Our system’s object independence assumption fails in cluttered environments, where object states and observations are influenced by nearby objects. This requires modifying our object-oriented belief update to handle these interactions. HOO-POMDP also cannot handle an unknown class of objects. While it is easy enough to categorize all such objects into a single ‘unknown’ class, the difficult part is to plan to find an empty space to move the unknown object to. In general, this could lead to complicated packing problems that are NP-hard.